OpenStudy (anonymous):

32^[(x+5)/(x-7)] = (.25)(128^[(x+17)/(x+3)]

jimthompson5910 (jim_thompson5910):

32^[(x+5)/(x-7)] = (.25)(128^[(x+17)/(x+3)] 32^[(x+5)/(x-7)] = (1/4)(128^[(x+17)/(x+3)] (2^5)^[(x+5)/(x-7)] = 2^(-2)((2^7)^[(x+17)/(x+3)] 2^[5(x+5)/(x-7)] = 2^(-2)(2^[7(x+17)/(x+3)] 2^[5(x+5)/(x-7)] = 2^[-2+(7(x+17))/(x+3)] Since the bases are equal (to 2), the exponents must be equal, so 5(x+5)/(x-7) = -2+(7(x+17))/(x+3) Solve that for x to find your answer

OpenStudy (anonymous):

I got that far. The answer is supposed to be 10, but I didn't get 10. I got some weird fraction. Is it right to first combine the terms on the right by finding the common denominator of x + 3?

jimthompson5910 (jim_thompson5910):

hmm I'm getting a fraction too, the original equation is this right \[\Large 32^{\frac{x+5}{x-7}} = (0.25)\left( 128^{\frac{x+17}{x+3}}\right ) \] ???

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

what fraction did you get for the answer?

OpenStudy (anonymous):

I don't remember but it was something about dividing 68 into something or something like that. I erased what I wrote

jimthompson5910 (jim_thompson5910):

I guess it doesn't matter, but I'm getting 433/19 for the answer. Not sure how they got 10.

OpenStudy (anonymous):

let me check 10

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

it doesn't check i don't think...

jimthompson5910 (jim_thompson5910):

hmm must be a typo somewhere, who knows

OpenStudy (anonymous):

probably, thanks anyway

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

can i message you later if I need help with another problem?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

[(x+17)/(x+3)] (x+17)/(x-3)

jimthompson5910 (jim_thompson5910):

hmm seems too easy, thanks Algebraic!

jimthompson5910 (jim_thompson5910):

32^[(x+5)/(x-7)] = (.25)(128^[(x+17)/(x-3)] 32^[(x+5)/(x-7)] = (1/4)(128^[(x+17)/(x-3)] (2^5)^[(x+5)/(x-7)] = 2^(-2)((2^7)^[(x+17)/(x-3)] 2^[5(x+5)/(x-7)] = 2^(-2)(2^[7(x+17)/(x-3)] 2^[5(x+5)/(x-7)] = 2^[-2+(7(x+17))/(x-3)] Since the bases are equal (to 2), the exponents must be equal, so 5(x+5)/(x-7) = -2+(7(x+17))/(x-3) Solve that for x to find your answer